OpenStudy (anonymous):

How do I find all possible values for Ɵ in the equation cos 4Ɵ = 0 ?

OpenStudy (anonymous):

cosx = 0 implies x = (2n+1) * pie/2 ............... hope nw u can find for cos 4x

OpenStudy (anonymous):

if i know that cos θ = 0 at \[\frac{ \pi }{ 2 } and \frac{ 3 \pi }{ 2 }\] then can i use this information to get the answer @Yahoo! ?

OpenStudy (larseighner):

What if those angles were 1/4 as big?

OpenStudy (anonymous):

so i should multiply those by 1/4 and those are the values for that equation?

OpenStudy (larseighner):

Well, if 4 times ? = one of those angles, wouldn't it follow that 4 times ? would satisify cos 4 theta = 0?

OpenStudy (larseighner):

Of course you realize you cannot really find ALL the valuse because cos theta is periodic, so the value recur as you go around the unit circle over and over.

OpenStudy (anonymous):

well when i enter cos 4 θ = 0 into mathway.com, i get \[\frac{ \pi }{ 8 } \pm \frac{ \pi n }{ 2 } , \frac{3 \pi}{2} \pm \frac{\pi n}{2}\] I'm not sure where the πn comes from.

OpenStudy (larseighner):

Okay, I misspoke there. If you know the values that satisfy \( \cos\theta = 0 \) then sure 1/4 of those values would satisfy \( \cos 4 \theta = 0 \). But you seem to have got it anyway.

OpenStudy (larseighner):

The unit circle is \( 2\pi \). Every time you go around it you get the same value. so if x is a solution, \(2\pi + x \) is a solution and so is \(4\pi + x \) and \(6\pi + x\) and so forth. So \(2n\pi + x\) is the a solution for integers n (except integer 0).

OpenStudy (anonymous):

I see, though why is it also divided by 2?

OpenStudy (larseighner):

I think that should be: \(\large\frac{ \pi }{ 8 } \pm \frac{ \pi n }{ 2 } , \frac{3 \pi}{8} \pm \frac{\pi n}{2}\) because of the plus/minus, it is telling you that it is true for n=1, 2, 3, ...

OpenStudy (anonymous):

ok, it makes more sense with it being as you just put than with 3π/2

OpenStudy (anonymous):

Thank you for the help.

OpenStudy (larseighner):

Sure.